From: Tony Cass
Sent: Thursday, June 22, 2006 6:56 AM
To: info-experiments (Master list for IT to broadcast to all experiment lists); it-dep-service-managers (IT Service Managers)

Subject: Services down after Computer Centre cooling problems

Dear all,

The temporary air conditioning equipment still needed to support the computer centre did not restart after a general 
powercut around 2am this morning. Non-critical IT services were shutdown as the temperature was rising rapidly and the 
machines powered off. Cooling was restored around 4am and we startd to resume services around 4:30. Unfortunately, the 
orderly shutdown procedure also included some database services in the critical area. Restoring physics services is 
therefore taking longer than expected.

More news around 10am.

Regards,
Tony 

-- JakubMoscicki - 22 Jun 2006

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-06-22 - JakubMoscicki
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    ArdaGrid All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback